Calculate the voltage of a area insulated susbstation if this outdoor substation has three busbars and supplied by units of post insulators. Each unit consists of 3 suspension type pin insulators with the voltage across the line unit is 9KV, while that across the middle unit is 7.5KV.
Answer #1

tur е HTT + HE مي ta K- al c. = 7.5kV We know, v=v Sol VL = Line voltage = Subotation Voltage V=9kr 3 = V(1tk) V4=V(1 +3K+) 9